Find the maximum possible domain and corresponding range of the function f defined by f(x) = {√(1–x²)/(x–2)}
1
Expert's answer
2019-10-22T12:23:09-0400


ANSWER: Domain of f is the set   "(-\infty,-1]\bigcup [1,2)" ; range of f is the set "[0,+\infty)"

EXPLANATION. If "x>2" , then the denominator of the fraction is positive and the numerator is negative. Therefore the fraction "\frac{1-x^{2}}{x-2}<0" . If "x<2" , then the domain of the function f consists of x, for which "1-{ x }^{ 2 }\le 0" , hence "x\in\left( - \infty,-1 \right ] \bigcup\left[1,+\infty \right )" . Thus, the domain of the function is the set "\left \{ (-\infty,-1]\bigcup [1,+\infty) \right \} \bigcap (-\infty,2) =" "(-\infty,-1]\bigcup [1,2)"

The range of the function is the set "[0,+\infty)" , since for any "a\geqslant0" the equation f(x)=a has a solution ,that belongs to [1,2). This follows from equivalent equalities :"x\in [1,2)\quad ,\sqrt { \frac { 1-{ x }^{ 2 } }{ x-2 } } =a\quad \Leftrightarrow \quad \frac { 1-{ x }^{ 2 } }{ x-2 } ={ a }^{ 2 }\quad \Leftrightarrow { \quad x }^{ 2 }+{ a }^{ 2 }x-2{ a }^{ 2 }-1=0\" "\Rightarrow x=\frac { -{ a }^{ 2 } }{ 2 } +\sqrt { \frac { { a }^{ 4 } }{ 4 } +2{ a }^{ 2 }+1 } \quad"